OpenStudy (anonymous):

An air-track glider of mass 0.25 kg moving at 0.60 m/s collides with and sticks to a glider of mass 0.50 kg at rest. How much kinetic energy is converted into other forms of energy as a direct result of this collision? A. 0.030 J B. 0.045 J C. 0.015 J D. 0.060 J ***not sure at all! :( Thank you!! :D

OpenStudy (michele_laino):

here we have to apply the conservation of momentum, so we can write: \[\Large {m_1}{v_1} = \left( {{m_1} + {m_2}} \right)u\]

OpenStudy (anonymous):

okay!

OpenStudy (michele_laino):

where m_1= 0.25 Kg, m_2=0.50 Kg, and v_1=0.60 m/sec

OpenStudy (michele_laino):

u is the final speed of the system particle1+ particle2

OpenStudy (michele_laino):

we have to find what is u: \[\Large u = \frac{{{m_1}{v_1}}}{{{m_1} + {m_2}}} = ...m/\sec \]

OpenStudy (anonymous):

okay, so i plug in like this? (0.25*0.60)=(0.25+0.50)u so like this? 0.15=0.75u u=0.2 ?

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

oh wait... it is 0.2, correct? the value of u?

OpenStudy (michele_laino):

that's right u=0.2 m/sec

OpenStudy (anonymous):

okay! what happens next?

OpenStudy (michele_laino):

Now the final kinetic energy is: \[\Large {E_f} = \frac{1}{2}\left( {{m_1} + {m_2}} \right){u^2} = ...Joules\]

OpenStudy (anonymous):

okay! so 1/2(0.25+0.50)0.2^2 = 0.015 J ? so the answer is 0.015 J? :O

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay!! thank you:D

OpenStudy (michele_laino):

whereas the initial kinetic energy is: \[\Large {E_i} = \frac{1}{2}{m_1}v_1^2 = ...Joules\]

OpenStudy (michele_laino):

@iheartfood

OpenStudy (anonymous):

ahh so it would be 1/2(0.25+0.5^2)? so 0.25 is the initial kinetic energy?

OpenStudy (michele_laino):

I got 0.045 Joules

OpenStudy (michele_laino):

\[\Large {E_i} = \frac{1}{2}{m_1}v_1^2 = \frac{1}{2} \times 0.25 \times {\left( {0.6} \right)^2} = ...Joules\]

OpenStudy (anonymous):

ohhh yes 0.045 sorry, i plugged in the wrong thing! wait so for this problem, i now have to do this to find the answer? 0.045-0.015=0.030 J ? and that would be this problem's final solution? :O

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

yay!! and so this problem is completed now? :O

OpenStudy (michele_laino):

yes! We have finished!

OpenStudy (anonymous):

yay!! thank you:D

OpenStudy (michele_laino):

thank you!! :D